OpenStudy (anonymous):

Evaluate the function for x = –2a, if a = 3. f(x) = –4x + 7

OpenStudy (science0229):

If a=3, x would be 3*-2, or -6. So plug -6 for x and find f(x). f(x)=(-4)(-6) + 7=24+7=31